# fifth generation 6G sixth generation

### v2

@@ IV Empirical Validation and Analysis
IV-A IFTR channel fitting
Once the Tx-Rx configurations have been selected, a fitting between the empirical PDFs and the IFTR theoretical one is carried out. A common way is the use of optimization numerical methods such as gradient descent algorithms or genetic algorithms (GA) to minimize an objective function. In this work, a multi-objective optimization problem has been defined with the following set of error metrics [^5] The $f_{\rm exp}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}(x)}$ and $f_{\rm mod}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}(x)}$ functions represent the empirical and the IFTR channel model PDFs respectively. $N$ corresponds with the number of points for representing the PDFs and is equal to 100 in all the studied configurations. as targets:
  • Mean Squared Error:
  • Root Mean Squared Error:
  • Mean Absolute Error:
  • Modified Kolmogorov-Smirnov statistic:
The use of multiple objective functions aims to find differences between the results when minimizing each error metric.
To solve the optimization problem, the chosen algorithm is the NSGA-II [31], an elitist GA designed for working with real numbers. It has been configured with a population size of 200 and a number of generations equal to 400. The elite count parameter (number of individuals guaranteed to survive to the next generation) has been established to the 5% of the population size. The algorithm has been configured to stop when 100 generations occur consecutively without significant changes in the solutions, which means they have converged. In addition, we have defined upper and lower bounds for limiting the search inside the solutions space. For the $K$ factor, we have restricted the search to the interval from 0 to 103. On the other hand, for the $m_{1}$ and $m_{2}$ parameters we have set 0 and 50 as upper and lower limits respectively. Finally, since the $\Delta$ parameter takes values in the interval [0,1], the search has been bounded to the complete definition domain.
For each fitted channel, the output of the GA is a set of $N_{p}$ solutions known as Pareto front. Solutions consist of a tuple of the form $\left(K,\Delta,m_{1},m_{2}\right)$ representing one compromise situation between the four objective functions. For instance, Fig. 4 shows the solutions obtained for the Pareto front of configuration A (see Fig. 2) in the anechoic case. Boxplot diagrams and swarm-chart scatter plots are represented jointly for each parameter (normalized with respect to its upper bound). The $K$ factor is distributed along a large interval which ranges from 0.36 to 0.97 (or equivalently from 25.6 dB to 29.8 dB, units no-normalized). However, most of the solutions are distributed near the median value of 0.50 (27.0 dB). In contrast, the $\Delta$ parameter practically does not change taking the same value of 0.44 for all the solutions tuples. Differences appear for the fluctuation parameters $m_{1}$ and $m_{2}$, emerging two well-defined values. Specifically, for the parameter $m_{1}$ are obtained 0.69 and 0.83 (34.5 and 41.5 in no-normalized units); meanwhile for the $m_{2}$ are obtained 0.32 and 0.46 (16 and 23 without normalizing). This particular behavior is common for the majority of the studied merged channel configurations in the three scenarios.
We can focus our attention on the solutions tied to the minimum of each error metric (see Table I). The two extreme values for the pair $m_{1}$ and $m_{2}$ are either associated to the average error metrics (MSE, RMS and MAE), or to the distance error metric (KS), respectively. As previously mentioned, $\Delta$ remains almost unchanged in the four solutions tuples. On the other hand, the $K$ factor is equal for the MSE and the RMSE solutions and differs roughly 1 dB with those obtained for the MAE and KS solutions. Fig. 5 shows the experimental PDF of the anechoic configuration A and the theoretical IFTR curves for each of the mentioned solutions, which are practically overlapped between each other. On the one hand, the PDFs associated with the MSE, RMSE and MAE error metrics have the same shape (quasi-equal IFTR parameters). On the other hand, the curve of the minimum KS is slightly different improving the fitting on the modal regions at the expense of the distribution tails. The latter are fitted with more accuracy by minimizing the mean errors.
To choose a particular solution $\hat{S}$ = $\left(K,\Delta,m_{1},m_{2}\right)$ from the Pareto front, we need some unified criterion. For that purpose, we first define the normalized error functions. As an example, the normalized $\mathrm{MSE}_{n}$ function is computed as
$\mathrm{MSE}_{n}(S)=\frac{\mathrm{MSE}(S)}{\displaystyle\max_{P}\mathrm{MSE}(P)}.$ (13)
$\varepsilon_{n}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}S})\triangleq\left.\frac{\mathrm{MSE}_{n}+\mathrm{RMSE}_{n}+\mathrm{MAE}_{n}+\mathrm{KS}_{n}}{4}\right|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}S}}.$ (14)
$\hat{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}S}}=\min_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}S}}\varepsilon_{n}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}S}).$ (15)
@@ IV-A IFTR channel fitting
The cumulative density functions of the optimal IFTR parameters for each scenario have been represented in Fig. 6. A high contrast between anechoic and multipath [^6] The term multipath refers, henceforth, to both reverberation and indoor propagation scenarios. scenarios is observed, with clear different trends.
Figure 6: Comparison of the optimal IFTR parameters CDFs for the three analysed environments.
Figure 7: RMSE metric CDFs for the optimal IFTR solutions.
TABLE I: IFTR parameters for each minimum error metric solution of the Pareto front in anechoic configuration A
Min. Error Metric Solutions (Conf. A)
Parameter MSE RMSE MAE KS
Min. error $16.56$ $4.07$ $2.96$ $8.18$
$K$ [$\mathrm{dB}$] $28.6$ $28.6$ $27.2$ $27.3$
$\Delta$ $0.44$ $0.44$ $0.44$ $0.44$
Fig. 6 shows that the Rician-like $K$ factor takes higher values on the anechoic configurations, which is due to the absence of diffuse power. In fact, the median value for this scenario is 28.3 dB, whereas for the multipath ones is 24.7 dB; that is, 3.6 dB of difference. On the other hand, as can be observed in Fig. 6(b), the $\Delta$ parameter in the anechoic case distributes along lower values than for the multipath configurations. That is a consequence of the lack of reflections in the anechoic scenario, where only the two specular rays are relevant and their amplitudes are significantly different. Moreover, it is also remarkable that anechoic and reverberation scenarios are the inverse of each other. For the latter, the multipath specular components that arrive at the antenna have similar amplitudes. That, in addition to the nLoS situation, imposes $\Delta$ taking higher values. The obtained median values are 0.30 for the anechoic case and 0.69 for reverberation and indoor scenarios. Although these two latter take the same median value, slight differences are observed between their distributions. For $\Delta$ values lower than the median, the distribution for the indoor configurations is higher than for the reverberation ones. In contrast, for higher $\Delta$ values, the roles are exchanged and the distribution becomes lower for the indoor case. This is explained by the mixed character of these types of scenarios. The studied indoor measurements combine the LoS situation with a rich multipath. The high power of the LoS rays produces an imbalance between the amplitudes of the received rays which implies a reduction of the $\Delta$ parameter.
The $m_{1}$ and $m_{2}$ fluctuation parameters distributions are depicted in Figs. 6(c) and 6(d). The first ray is more fluctuating in the reverberation and indoor scenarios than in the anechoic ones since the $m_{1}$ parameter is distributed along low values (the medians are 1 and 28 respectively). That is due to the multipath effect, which provides more accentuated variations in the received signal. This also explains that distribution for reverberation and indoor are overlapped between each other. On the other hand, the $m_{2}$ parameter takes higher values than the $m_{1}$ in the multipath scenarios (median values of 20 and 21 respectively). This implies that, generally, the second ray is less fluctuating than the first for the reverberation and indoor cases, whereas for the anechoic both rays have reduced fluctuation. Nevertheless, for the indoor configurations, there exists a remarkable percentage of channels with a low $m_{2}$ parameter which is related to a significant multipath effect.
@@
TABLE II: Numerical results of the given fitting example configurations in the three studied scenarios
@@
Figure 8: PDF fitting examples for different configurations. Solid and dotted lines represent empirical and IFTR model distributions respectively.
To compare the fitting accuracy between the analyzed scenarios, the RMSE metric associated with each solution has been studied. Fig. 7 shows the CDFs for the three scenarios. The lowest error is obtained for the indoor case, with a median value of 3.2; whereas for the reverberation and anechoic cases the medians are 7.0 and 23.0 respectively. From this, it can be affirmed that the IFTR model describes properly rich multipath scenarios in both LoS and nLoS situations. The largest RMSE values correspond to the anechoic case. As it will be discussed later, this is due to a non-uniform phase distribution in the received signal providing some contradictions with the IFTR model underlying assumptions.
Some particular fitting examples [^7] As supplementary material for the reader, an additional animation is included showing the obtained fitting solutions for each of the 142 configurations in the three studied scenarios. configurations (namely as A, B, C, D, E, and F) are depicted in Fig. 8. A good fitting accuracy is observed for the three scenarios, and the particular numerical results can be consulted in Table II. Bimodal behaviors are accurately modeled, as well as Rician- and Rayleigh-like ones common in multipath scenarios. Specifically, the IFTR model is capable of reproducing bimodality as long as $K$ takes large values and $\Delta$ is near to one. These conditions are equivalent to those explained in Section III for experimental bimodality. That is, the power of the first ray has to be comparable to the one of the second ray, and both have to be notably higher than the diffuse one.
@@
Fig. 9 shows the specific configurations of the given examples in terms of the two combined Rx array positions (points with the same color) jointly with the common Tx orientation. Positions are referred with respect to the array center, which corresponds to the (0,0) coordinate.
@@ IV-B Anechoic environment analysis
As previously mentioned, anechoic scenarios are associated with higher fitting error than multipath ones (see Fig. 7). Although most of the configurations can be described with high accuracy, there exists a certain percentage of fitted configurations with an unsatisfactory goodness of fit (GoF). After a thorough analysis, the key aspect that justifies this behavior is the lack of phase diversity in certain combinations. That is, the traveled path by the waves and the deep attenuation of diffuse components is insufficient to produce all phase differences between $-\pi$ and $\pi$. This has two main implications: on the one hand, it causes the CLT assumption for the diffuse component is not met; on the other hand, it affects the assumption of uniformly distributed phases for the dominant specular components. The latter is of importance because you can write the received amplitude of the IFTR model in terms of the phase difference $\alpha$ as follows:
@@
The implications of a non-uniform phase distribution have been discussed in [21], where the TWDP model was modified to account for this particular situation. The resulting model, termed as generalized two-ray model with von Mises phase distribution (GTR-V), has the following PDF for the received amplitude $r$:
where $f_{\rm rice}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}r,K_{r}})$ is the well-known Rician distribution with a factor $K_{r}=K[1+\Delta\cos(\alpha)]$. The PDF is given by
Figure 10: Phase difference fitting with von Mises distribution for the anechoic configuration G.
Indeed, it is possible to generalize the IFTR model to incorporate the effect of a non-uniform phase distribution following the rationale in [21]. Nevertheless, this would imply that the resulting model would have two additional parameters, incurring additional complexity when it comes to fitting. Instead, we will use the GTR-V model even though it has no fluctuation on the two dominant specular components; in practice, this is equivalent to assuming a sufficiently large value of $m_{1}$ and $m_{2}$ in the IFTR model (see Fig. 6). Therefore, the use of this alternative model is well-justified for the sake of simplicity. After analyzing the distribution of the phase difference between the dominant specular components in the relevant anechoic configurations, a small sub-set with von Mises-like PDF for phase difference has been selected (named configurations G, H, and I).
(a) Configuration G.
(b) Configuration H.
(c) Configuration I.
Figure 11: Comparison between fitted IFTR and GTR-V PDFs for some particular anechoic configurations.
Fig. 10 shows the fitting of the phase difference distribution [^8] Though the channel measurements include both diffuse and specular components, in anechoic channels the LoS ray power is significantly the highest. Thus, it is reasonable to estimate the phase difference distribution as the PDF of the phase difference of the two merged channels. with a von Mises PDF for configuration G. After the first step of the optimization, the resulting parameters $\kappa$ and $\phi$ yield an MSE of 0.0029, which implies an excellent GoF. This process has been carried out also for H and I configurations. The next step is to find the optimal $K$ and $\Delta$ parameters for the GTR-V model by the use of GA. The obtained results for the configurations G, H, and I are summarized in Table III, comparing them with the obtained for the IFTR model. In addition, the obtained PDFs are depicted in Fig. 11. For configurations G and H [Figs. 11(a) and 11(b)], we see that the GTR-V model provides better results than the IFTR one. The effect of a non-uniform phase distribution is translated into an effective reduction of the variance, and it can be visually confirmed that the GTR-V model achieves this. However, we see that the fitting is not improved when the empirical PDF is bimodal – see configuration I in Fig. 11(c). The capabilities of the TWDP model to capture bimodality are affected when a non-uniform phase difference distribution is considered in the GTR-V model.
TABLE III: Numerical results comparison between the IFTR and the GTR-V models
@@
Figure 12: PDFs empirical and fitted for sharp bimodality behaviors.
@@ IV-C Multipath environments analysis
Multipath environments have lower RMSEs compared to the anechoic ones, as has been shown in Fig. 7. This implies that most of the configurations have been described correctly with the IFTR model. However, we have identified some cases in which the fitting fails. Specifically, in the cases of sharp bimodal behaviors, the ability of the IFTR model to capture such bimodality is not enough. This is exemplified in Fig. 12 for another configurations J and K in reverberation and indoor environments respectively. In these situations, the model tries to fit the dominant peak but masks the secondary one. Numerical results can be consulted in Table IV.
TABLE IV: IFTR numerical results obtained for multipath configurations with sharp bimodality behaviors
@@
(a) Configuration J (reverberation).
(b) Configuration K (indoor).
Figure 13: CIR for multipath configurations with sharp bimodality behaviors.
One way to determine the origin of these particular behaviors is to analyze the CIR of the configuration. Fig. 13 shows the estimation of the channel time responses. On the one hand, Fig. 13(a) depicts the CIR for the reverberation configuration J. It is clear that more than two rays are arriving at the antenna with power significantly high and comparable between each other. This is not covered by the IFTR model. Even though the consideration of additional rays can be usually encompassed by integrating these into the diffuse component [30], this is not the case when the overall number of rays is reduced, multipath propagation is limited, and there is a lack of phase richness in the propagation environment. On the other hand, Fig. 13(b) presents the CIR for the indoor configuration J. Two dominant rays are observed, but they are overlapped in time between each other. Hence, its interaction may not be described independently as the IFTR model assumes. The sharp bimodality behavior appears as a cause of the time interaction of these two rays.
@@ V Conclusions
We conducted an empirical validation of the independent fluctuating two-ray (IFTR) fading model, covering a wide range of scenarios in the frequency band n258 (from 24.25 GHz to 27.5 GHz). We confirmed that the IFTR model is highly versatile, and is able to recreate rather dissimilar propagation conditions including anechoic, reverberation, and indoor scenarios. Besides, the four shape parameters that characterize the model have a solid physical meaning, which agrees with the expected properties of each analyzed channel.
@@
We also observed some relevant effects that put forth the limitations of the IFTR model: (i) the assumption of uniform phases for the dominant specular components does not always hold, especially when there is a lack of phase richness. This is the case in scenarios with a reduced multipath, and is likely to become a dominant effect as we move up in frequency; (ii) the IFTR model is not able to capture sharp bimodal behaviors observed in some of the measurements. This can be caused by the combination of a number of effects, such as the presence of additional rays (due to spurious reflections), the absence of rich multipath propagation, and the interaction between dominant specular components. In this regard, the development of physically-motivated models that capture these behaviors deserves special attention, together with additional empirical validations at higher frequencies.
Appendix

### v3

@@ IV Empirical Validation and Analysis
IV-A Analysis of the obtained IFTR parameters
@@
Once the Tx-Rx configurations have been selected, a fitting between the empirical PDFs and the IFTR theoretical one is carried out. In this work, we have decided to use a methodology based on multi-objective optimization with a genetic algorithm (GA). For a detailed explanation and certain particularities on the fitting methodology, see Appendix A. The cumulative density functions of the obtained IFTR parameters for each scenario are represented in Fig. 4. A high contrast between anechoic and multipath [^6] The term “multipath” refers, henceforth, to both reverberation and indoor propagation scenarios. scenarios is observed, with clear different trends.
@@
TABLE I: Numerical results obtained for the given fitting examples in the three studied scenarios
@@
Figure 5: Fitting examples for different configurations in the three analysed scenarios. Solid and dotted lines represent empirical and IFTR model distributions respectively.
@@
From Fig. 4, it can be seen that the $K$-factor takes higher values on the anechoic configurations, which is due to the absence of diffuse power. In fact, the median value for this scenario is 28.3 dB, whereas for the multipath ones is 24.7 dB; that is, 3.6 dB of difference. The graphic also shows the maximum measurable $K$-factor for an individual ray at a distance of 1 meter for the given VNA setup. It has been empirically determined in our anechoic chamber, and takes values up to 27.5 dB. The limit for two-rays would be 3 dB higher (at most), and since the distances between Tx and Rx in our measurements are significantly larger than 1 m, we can ensure that there is no merged channel which might surpass an IFTR $K$-factor of 30 dB as confirmed in Fig. 4.
On the other hand, as can be observed in Fig. 4(b), the $\Delta$ parameter in the anechoic case distributes along lower values than for the multipath configurations. That is a consequence of the lack of reflections in the anechoic scenario, where only the two specular rays are relevant and their amplitudes are significantly different. Moreover, it is also remarkable that anechoic and reverberation scenarios are the inverse of each other. For the latter, the multipath specular components that arrive at the antenna have similar amplitudes. That, in addition to the nLoS situation, imposes $\Delta$ taking higher values. The obtained median values are 0.30 for the anechoic case and 0.69 for reverberation and indoor scenarios. Although these two latter take the same median value, slight differences are observed between their distributions. For $\Delta$ values lower than the median, the distribution for the indoor configurations is higher than for the reverberation ones. In contrast, for higher $\Delta$ values, the roles are exchanged and the distribution becomes lower for the indoor case. This is explained by the mixed character of these types of scenarios. The studied indoor measurements combine the LoS situation with a rich multipath. The high power of the LoS rays produces an imbalance between the amplitudes of the received rays which implies a reduction of the $\Delta$ parameter.
The $m_{1}$ and $m_{2}$ fluctuation parameters distributions are depicted in Figs. 4(c) and 4(d). The first ray fluctuates more in the reverberation and indoor scenarios than in the anechoic ones since the $m_{1}$ parameter is distributed along low values (the medians are 1 and 28 respectively). That is due to the multipath effect, which provides more accentuated variations in the received signal. This also explains that distribution for reverberation and indoor are overlapped between each other. On the other hand, the $m_{2}$ parameter takes higher values than the $m_{1}$ in the multipath scenarios (median values of 20 and 21 respectively). This implies that, generally, the second ray is less fluctuating than the first for the reverberation and indoor cases, whereas for the anechoic both rays have reduced fluctuation. Nevertheless, for the indoor configurations, there exists a remarkable percentage of channels with a low $m_{2}$ parameter which is related to a significant multipath effect.
Some particular fitting examples [^7] As supplementary material for the reader, an additional animation is included showing the obtained fitting solutions for each of the 142 configurations in the three studied scenarios (see Appendix B). configurations (namely as A, B, C, D, E, and F) are depicted in Fig. 5. A good fitting accuracy is observed for the three scenarios, and the particular numerical results can be consulted in Table I. Bimodal behaviors are accurately modeled, as well as Rician-like and Rayleigh-like ones common in multipath scenarios. Specifically, the IFTR model is capable of reproducing bimodality as long as $K$ takes large values and $\Delta$ is near to one. These conditions are equivalent to those explained in Section III for experimental bimodality. That is, the power of the first ray has to be comparable to the one of the second ray, and both have to be notably higher than the diffuse one. Finally, Fig. 6 shows the specific configurations of the given examples in terms of the two combined Rx array positions (points with the same color) jointly with the common Tx orientation. Positions are referred with respect to the array center, which corresponds to the (0,0) coordinate.
The IFTR model is able to properly describe the expected characteristics from such diverse scenarios (anechoic, reverberation and indoor environments) under very different conditions. Since from the set of merged channels multiple behaviors emerge, this ray-based model exhibits a great versatility for describing complex scenarios. As mentioned in the previous section, these merged channels may resemble a great variety of indoor scenarios (e.g. offices, meeting rooms…). Thus, the model applicability covers multitude of different channels and environments.
@@ IV-B Anechoic environment analysis
Although most of the anechoic configurations can be described with high accuracy, there exists a certain percentage of fitted configurations with an unsatisfactory GoF. After a thorough analysis, the key aspect that justifies this behavior is the lack of phase diversity in certain combinations. That is, the traveled path by the waves and the deep attenuation of diffuse components is insufficient to produce an uniform distribution of phase differences between $-\pi$ and $\pi$. This has two main implications: on the one hand, it causes the CLT assumption for the diffuse component is not met; on the other hand, it affects the assumption of uniformly distributed phases for the dominant specular components. The latter is of importance since the received amplitude of the IFTR amplitude can be formulated in terms of the phase difference $\alpha$ as follows:
@@
Figure 7: Phase difference distribution for the anechoic configuration G. The dashed line represents the fit with the von Mises PDF.
TABLE II: Numerical results comparison between the IFTR and the GTR-V models
@@
(a) Configuration G.
(b) Configuration H.
(c) Configuration I.
Figure 8: Comparison between fitted IFTR and GTR-V PDFs for several anechoic configurations.
The implications of a non-uniform phase distribution have been discussed in [23], where the TWDP model was modified to account for this particular situation. The resulting model, termed as GTR-V, has the following PDF for the received amplitude $r$:
where $f_{\rm rice}({r,K_{r}})$ is the well-known Rician distribution with a factor $K_{r}=K[1+\Delta\cos(\alpha)]$. The PDF is given by
Indeed, it is possible to generalize the IFTR model to incorporate the effect of a non-uniform phase distribution following the reasoning of [23]. Nevertheless, this would imply that the resulting model would have two additional parameters, incurring additional complexity when it comes to fitting. Instead, we will use the GTR-V model even though it has no fluctuation on the two dominant specular components; in practice, this is equivalent to assuming a sufficiently large value of $m_{1}$ and $m_{2}$ in the IFTR model (see Fig. 4). Therefore, the use of this alternative model is well-justified for the sake of simplicity. After analyzing the distribution of the phase difference between the dominant specular components in the relevant anechoic configurations, a small sub-set with von Mises-like PDF for phase difference has been selected (named configurations G, H, and I).
@@
Figure 9: Empirical and fitted PDFs for sharp bimodality behaviors.
Fig. 7 shows the fitting of the phase difference distribution [^8] Though the channel measurements include both diffuse and specular components, in anechoic channels the LoS ray power is significantly the highest. Thus, it is reasonable to estimate the phase difference distribution as the PDF of the phase difference of the two merged channels. with a von Mises PDF for configuration G. After the first step of the optimization, the resulting parameters $\kappa$ and $\phi$ yield an MSE of 0.0029, which implies an excellent GoF. This process has been carried out also for H and I configurations. The next step is to find the optimal $K$ and $\Delta$ parameters for the GTR-V model by the use of GA. The obtained results for the configurations G, H, and I are summarized in Table II, comparing them with the obtained for the IFTR model. In addition, the obtained PDFs are depicted in Fig. 8. For configurations G and H [Figs. 8(a) and 8(b)], it is observed that the GTR-V model provides better results than the IFTR model. The effect of a non-uniform phase distribution is translated into an effective reduction of the variance, and it can be visually confirmed that the GTR-V model achieves this. However, we see that the fitting is not improved when the empirical PDF is bimodal — see configuration I in Fig. 8(c). Thus, the capabilities of the TWDP model to capture bimodality are affected due to the non-uniform phase difference distribution.
(a) Configuration J (reverberation).
(b) Configuration K (indoor).
Figure 10: CIRs for multipath configurations with sharp bimodality behaviors.
@@ IV-C Multipath environments analysis
For the multipath cases, most of the configurations are described correctly with the IFTR model. However, we have identified some cases where it fails. Specifically, in situations with sharp bimodal behaviors, the ability of the IFTR model to capture such bimodality is not enough. This is exemplified in Fig. 9 for another configurations J and K in reverberation and indoor environments respectively. In these situations, the model tries to fit the dominant peak but masks the secondary one. Numerical results can be consulted in Table III.
TABLE III: Numerical results of the fitting between the IFTR model and the multipath channels with sharp bimodality behaviors
@@
One way to determine the origin of these particular behaviors is to analyze the CIR module of the configurations (see Fig. 10). On the one hand, Fig. 10(a) depicts the CIR for the reverberation configuration J. It is clear that more than two rays are arriving at the antenna with power significantly high and comparable between each other. This is not covered by the IFTR model. Even though the consideration of additional rays can be usually encompassed by integrating these into the diffuse component [32], this is not the case when the overall number of rays is reduced, multipath propagation is limited, and there is a lack of phase richness in the propagation environment. On the other hand, Fig. 10(b) presents the CIR for the indoor configuration K. Two dominant rays are observed, but they are overlapped in time between each other. Hence, its interaction may not be described independently as the IFTR model assumes. The sharp bimodality behavior appears as a cause of the time interaction of these two rays.
@@ V Conclusions
We conducted an empirical validation of the independent fluctuating two-ray (IFTR) fading model, covering a wide range of scenarios in the frequency band n258 (from 24.25 GHz to 27.5 GHz). We confirmed that the IFTR model is highly versatile, and is able to recreate rather dissimilar propagation conditions including anechoic, reverberation, and indoor scenarios. Besides, the four shape parameters that characterize the model have a solid physical meaning, which agrees with the expected properties of each analyzed channel. That is, these few parameters are sufficient to successfully capture the particularities of propagation environments without the need to resort to more complex models with a larger number of parameters.
@@
We also observed some relevant effects that put forth the limitations of the IFTR model: (i) the assumption of uniform phases for the dominant specular components does not always hold, especially when there is a lack of phase richness. This is the case in scenarios with a reduced multipath, and is likely to become a dominant effect as we move up in frequency; (ii) the IFTR model is not able to capture sharp bimodal behaviors observed in some of the measurements. This can be caused by the combination of a number of effects, such as the presence of additional rays (due to spurious reflections), the absence of rich multipath propagation, and the interaction between dominant specular components. In this regard, the development of physically-motivated models that capture these behaviors deserves special attention, together with additional empirical validations at higher and lower frequencies; and also in new different scenarios.
Appendix A
IFTR Channel Model Fitting
Although moment-based or maximum-likelihood (ML) estimators are often used with sample data to fit a model distribution, cases where the distribution has a complicated form often require to be solved numerically [36]. Alternatively, the empirical PDF can be estimated and then, an optimization problem is solved to obtain the values of the model parameters that minimize an error metric. In our case, to perform the fit between the IFTR model and our measurements we have applied a methodology based on GA [^9] Other alternatives such as the MATLAB in-built routines fminsearch and lsqcurvefit were also implemented [37], obtaining a comparable but slightly worse performance compared to the GA method [38]., specifically the so-called non-dominated sorting genetic algorithm II (NSGA-II) [39].
A multi-objective optimization problem has been defined with the following set of error metrics [^10] The $f_{\rm exp}(x)$ and $f_{\rm mod}(x)$ functions represent the empirical and the IFTR channel model PDFs respectively. The parameter $N$ = 100, and corresponds with the number of PDF sampled points. as targets:
  • Mean Squared Error (MSE):
  • Root Mean Squared Error (RMSE):
  • Mean Absolute Error (MAE):
  • Modified Kolmogorov-Smirnov (KS) statistic:
The use of multiple objective functions provides the GA of a greater robustness to search for optimal solutions. Additionally, it aims to find differences between the results when minimizing each error metric.
As an evolutionary algorithm, NSGA-II requires both a population size and a maximum number of generations to be defined. In our specific case, since we only need to tune four parameters, a population of 200 individuals and 400 generations is adequate. Another relevant parameter is the elite count (number of individuals guaranteed to survive to the next generation), which has been established to the 5% of the population size. Finally, as stopping condition, we have configured the algorithm to stop when 100 generations occur consecutively without significant changes in the solutions.
Figure 11: RMSE metric CDFs for the optimal IFTR solutions.
Another relevant concept in the optimization problem is the solutions space, which is infinite. Thus, it is necessary to define upper and lower bounds for limiting the search along each variable. For the $K$-factor, we have restricted the search to the interval from 0 to 103 (that is, up to 30 dB). On the other hand, for the $m_{1}$ and $m_{2}$ parameters we have set 0 and 50 as upper and lower limits respectively. Finally, since the $\Delta$ parameter takes values in the interval [0,1], the search has been bounded to the complete definition domain. As can be seen in Fig. 4, there are no discontinuities on the distributions at the upper/lower boundary values. For instance, if it was not the case, it would mean that the GA saturated into the limit value of 30 dB in the case of the $K$-factor. Therefore, the established bounds for the GA search are adequate providing a good GoF as Fig. 5 shows.
For each fitted channel, the output of the GA is a set of $N_{p}$ solutions known as Pareto front. Each solution consists of a tuple $(K,\Delta,m_{1},m_{2})$ representing a compromise situation between the four objective functions. To choose a particular solution $\hat{S}$ = $(\hat{K},\hat{\Delta},\hat{m}_{1},\hat{m}_{2})$ from the Pareto front, we need some unified criterion. For that purpose, we first define the normalized error functions. As an example, the normalized $\mathrm{MSE}_{n}$ function would be
$\mathrm{MSE}_{n}(S){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\triangleq}\frac{\mathrm{MSE}(S)}{\displaystyle\max_{P}\mathrm{MSE}(P)}.$ (17)
$\varepsilon_{n}({S})\triangleq\left.\frac{\mathrm{MSE}_{n}+\mathrm{RMSE}_{n}+\mathrm{MAE}_{n}+\mathrm{KS}_{n}}{4}\right|_{{S}}.$ (18)
$\hat{{S}}=\min_{{S}}\varepsilon_{n}({S}).$ (19)
This equally-weighted normalized metric has been selected for the sake of simplicity. Since all the solutions from the Pareto front result in similar absolute error metrics, the choice of weights was not a key factor during our experiments. We give the distribution of the RMSE for the selected solutions as a reference for the fitting accuracy (see Fig. 11). The lowest error is obtained for the indoor case, with a value of 3.2; whereas for the reverberation and anechoic cases the medians are 7.0 and 23.0 respectively.
TABLE IV: IFTR parameters for each minimum error metric solution of the Pareto front in anechoic configuration A
Min. Error Metric Solutions (Conf. A)
Parameter MSE RMSE MAE KS
Min. error $16.56$ $4.07$ $2.96$ $8.18$
$K$ [$\mathrm{dB}$] $28.6$ $28.6$ $27.2$ $27.3$
$\Delta$ $0.44$ $0.44$ $0.44$ $0.44$
Although one solution from the complete Pareto front has to be selected, it is interesting to analyze the effects of the individual error functions on the fits. For instance, Fig. 12 shows the solutions obtained for the Pareto front of configuration A (see Fig. 2) in the anechoic case. Boxplot diagrams and swarm-chart scatter plots are represented jointly for each parameter (normalized with respect to its upper bound). The $K$-factor is distributed along a large interval which ranges from 0.36 to 0.97 (or equivalently from 25.6 dB to 29.8 dB, units no-normalized). However, most of the solutions are distributed near the median value of 0.50 (27.0 dB). In contrast, the $\Delta$ parameter practically does not change taking the same value of 0.44 for all the solutions tuples. Differences appear for the fluctuation parameters $m_{1}$ and $m_{2}$, emerging two well-defined values. Specifically, for the parameter $m_{1}$ are obtained 0.69 and 0.83 (34.5 and 41.5 in no-normalized units); meanwhile for the $m_{2}$ are obtained 0.32 and 0.46 (16 and 23 without normalizing). This particular behavior is common for the majority of the studied merged channel configurations in the three scenarios. We can focus our attention on the solutions tied to the minimum of each error metric (see Table IV). The two values for the pair $m_{1}$ and $m_{2}$ are either associated to the average error metrics (MSE, RMS and MAE) or to the distance error metric (KS) respectively. As previously mentioned, $\Delta$ remains almost unchanged in the four solutions tuples. Finally, the $K$-factor is equal for the MSE and the RMSE solutions and differs roughly 1 dB with those obtained for the MAE and KS solutions. In Fig. 13 are shown the experimental PDF of the anechoic configuration A and the theoretical IFTR curves for each of the mentioned solutions, which are practically overlapped between each other. On the one hand, the PDFs associated with the MSE, RMSE and MAE error metrics have the same shape (quasi-equal IFTR parameters). On the other hand, the curve of the minimum KS is slightly different improving the fitting on the modal regions at the expense of the distribution tails. The latter are fitted with more accuracy by minimizing the mean errors.
Appendix B
Supplementary Material